# A new twist on the Majorana surface code: Bosonic and fermionic defects for fault-tolerant quantum computation

### v3

@@ V Overheads for Universal Fault-Tolerant Computation
A Majorana triangle code is a patch of MSC encoding a single logical MZM. It has three odd-length boundaries of different colors (see Figure 16a). We begin by preparing $n$ magic states in $4n$ Majorana triangle codes, using for each quartet the “off-patch” method outlined in Section IV.4.2, followed by additional measurements to sever the “tails” in Figure 13b. (We assume having high-fidelity magic states, otherwise we can employ magic state distillation to improve the magic-state quality.) Let the distance of this encoding be $2d_{\Delta}$; that is, the minimum weights of the logical MZMs hosted on the four triangle codes are all $d_{\Delta}$. Meanwhile, let the distance of any bosonic MSC patch be $d$. We will assume we are in a parameter regime in which $d_{\Delta}<d$ and discuss the possibility of this below.
@@
Previous work demonstrated computation with Majorana triangle codes but utilized only bosonic aspects of the MSC for the ancilla patches containing magic states [Maj_Triangle_Code2018]. This yielded a scheme where magic state injection required an $O(d^{2})$ number of ancilla Majoranas in addition to four triangle codes: The ancilla cost is at least that of a distance-$d$ patch (e.g., that of Figure 1). This compares to the $O(1)$ number of ancilla Majoranas [namely the 4 MZMs used to prepare the noisy state for injection, and the $O(1)$-length tails] required in the off-patch method from Section IV.4.2. Thus, for large $d$, utilizing fermionic aspects of the MSC in magic state preparation and storage becomes more efficient. (Note that it is not possible to use bosonic ancilla patches for measurement as we require odd-length boundaries for lattice surgery with triangle codes; see Table 3.)
@@
To consider the overheads associated with fermionic ancilla patches for logical measurement, we first discuss error processes that we must contend with in order for our procedures to be fault-tolerant. Measurement errors may make thin-width sections of code vulnerable to unobservable, low-weight error mechanisms, depending on the patch preparation and readout used. We illustrate this error through the example shown in Figure 16b (black ring). In this example, ancilla patch preparation starts by preparing blue bond operators in their $+1$ eigenstates, followed by the measurement of blue plaquettes, which all have random outcomes. A fermion-parity-odd string that terminates on red and green boundaries can flip a blue plaquette, as shown in Figure 16b. However, if such an error string occurs shortly after state preparation along blue bonds, measurement errors can make it unobservable: suppose, for example, a blue plaquette suffers an initial measurement error, and it is then flipped by the parity-odd string described above. Subsequent measurements of the blue plaquette cannot detect the error string because of the initial measurement error (see Ref. Anyon_condensation_floquet_codes for a discussion of analogous error mechanisms in the bosonic Color Code).
@@
VI Fault-Tolerant Gates on Small Majorana Surface Codes
@@
In Figure VI, we show how braiding two $Z_{3}$ twists in a quartet affects the logical operators $\bar{X}_{1}$, $\bar{Z}_{2}$ and $\bar{Z}_{1}$ for the two BT qubits. The effect on logical operator $\bar{X}_{2}$ is shown in Figure 10 of the main text.
@@ Appendix E Lattice Surgery for Large-Weight Logical MZM Bilinear Measurement
In Figure E, we illustrate the lattice surgery procedure for measuring the bilinear operator $i\Upsilon_{a}\Upsilon_{b}$, where these two logical MZMs have arbitrarily high weights. In the example, $\Upsilon_{a}$ is supported on a blue boundary and $\Upsilon_{b}$ on a red boundary. Similarly to in Section IV.3.1, we prepare a patch, supporting ancilla logical MZMs $\tilde{\Upsilon}_{a}$ and $\tilde{\Upsilon}_{b}$, in the state $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}=+1$. These logical MZMs will have a similar weight to $\Upsilon_{a}$ and $\Upsilon_{b}$ respectively, and will be supported on blue boundaries. We measure the extra plaquette operators shown in Figure 23b connecting the two patches. Note that the product of the eigenvalues of those plaquette operators connecting $\Upsilon_{a}$ and $\tilde{\Upsilon}_{a}$ is equal to the eigenvalue of $i\Upsilon_{a}\tilde{\Upsilon}_{a}$, and similarly for $i\Upsilon_{b}\tilde{\Upsilon}_{b}$. We then measure $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}$ destructively (by measuring all blue bond operators) and perform/track a correction operation dependent on this measurement outcome (cf. Section IV.3.1).
@@ Appendix F Calculating the Correction Operation for Logical Fermion Parity Measurement
Here, we go through the details of Section IV.3.2, specifically showing how to determine the correction operation that needs to be tracked when measuring a logical fermion parity operator.
To determine this operation, we need to assess whether each logical bilinear operator has been preserved by the measurement-based procedure outlined. We deal with the example of $i\Upsilon_{1}\Upsilon_{3}$ specifically. We can use the final column of Table 2 to determine that after the lattice surgery measurements, this bilinear is mapped to $\lambda_{1}\lambda_{3}\bar{Z}_{B}$ (it is unchanged by the final measurement of $i\Upsilon_{4}\tilde{\Upsilon}_{4}$). The final step of the procedure from Section IV.3.2 is to measure all blue bond operators on the ancilla patch, from which we deduce the measurement outcome of operator $\tilde{\Gamma}_{4}$ to be $\eta_{1}$, $i\tilde{\Upsilon}_{1}\tilde{\Upsilon}_{2}\,\bar{X}_{B}$ to be $\eta_{2}$ and $i\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{4}\,\bar{Z}_{B}$ to be $\eta_{3}$ (these measurements do not commute with $\bar{Z}_{B}$). The post-measurement state of the logical qubits can be obtained by acting three logical operators $R_{1}$, $R_{2}$ and $R_{3}$ on the pre-measurement state. This is equivalent to replacing projectors with logical braids in Section IV.3.1.
We now determine $R_{1}$, $R_{2}$ and $R_{3}$ for this example. Note the logical operators stabilizing the pre-measurement state are those that were just measured via lattice surgery: $\lambda_{j}i\Upsilon_{j}\tilde{\Upsilon}_{j}$ for $j=1,2,3,4$. We consider first the measurement of $\tilde{\Gamma}_{4}$ (it does not matter which we choose to consider first, since the measurements all commute). A logical braid that enacts the same transformation as this measurement is (see Section IV.3.1):
@@ Appendix G Arbitrary String Operator Measurement Details
Here we extend the procedures of Section IV.5 to measuring arbitrary products of fermionic and bosonic string operators. We assume we wish to measure a product of $k$ operators, labeled $S_{1},\ldots,S_{k}$ and colored either red or blue. We start by introducing $k$ ancilla patches, each hosting two logical MZMs, $\tilde{\Upsilon}_{j_{a}}$ and $\tilde{\Upsilon}_{j_{b}}$ for $j=1,\ldots,k$, and we prepare each patch in the $i\tilde{\Upsilon}_{j_{a}}\tilde{\Upsilon}_{j_{b}}=+1$ state. This also prepares the $+1$ eigenstate of operator $\tilde{\Gamma}=\prod_{j=1}^{k}i\tilde{\Upsilon}_{j_{a}}\tilde{\Upsilon}_{j_{b}}$. An example of this setup is shown in Figure LABEL:subfig:arb_mmt_app_1 for $k=2$. In this example, $S_{1}$ is a bosonic string operator encircling a pair of fermionic twists hosting logical MZMs on blue boundaries. Both $S_{1}$ and $S_{2}$ are red string operators. We then measure $i\tilde{\Upsilon}_{j_{b}}\tilde{\Upsilon}_{(j+1)_{a}}$ for all $j<k$, to entangle these ancilla patches, while preserving the state of $\tilde{\Gamma}$. To achieve this, we can bridge the gap between ancilla patches with other ancilla patches. In Figure LABEL:subfig:arb_mmt_app_1, we introduce a patch hosting $\tilde{\Upsilon}_{c}$ and $\tilde{\Upsilon}_{d}$. The measurement of $i\tilde{\Upsilon}_{1_{b}}\tilde{\Upsilon}_{2_{a}}$ is performed via lattice surgery, shown in Figure 24b (see also Section IV.3.1 for an explanation of this procedure). After this measurement, the middle patch hosting $\tilde{\Upsilon}_{c}$ and $\tilde{\Upsilon}_{d}$ is discarded.

### v4

@@ 5 Overheads for Universal Fault-Tolerant Computation
A Majorana triangle code is a patch of MSC encoding a single logical MZM. It has three odd-length boundaries of different colors (see Figure 16a). We begin by preparing $n$ magic states in $4n$ Majorana triangle codes, using for each quartet the “off-patch” method outlined in Section 4.4.2, followed by additional measurements to sever the “tails” in Figure 13b. (We assume having high-fidelity magic states, otherwise we can employ magic state distillation to improve the magic-state quality.) Let the distance of this encoding be $2d_{\Delta}$; that is, the minimum weights of the logical MZMs hosted on the four triangle codes are all $d_{\Delta}$. Meanwhile, let the distance of any bosonic MSC patch be $d$. We will assume we are in a parameter regime in which $d_{\Delta}<d$ and discuss the possibility of this below.
@@
Previous work demonstrated computation with Majorana triangle codes but utilized only bosonic aspects of the MSC for the ancilla patches containing magic states [Maj_Triangle_Code2018]. This yielded a scheme where magic state injection required an $O(d^{2})$ number of ancilla Majoranas in addition to four triangle codes: The ancilla cost is at least that of a distance-$d$ patch (e.g., that of Figure 1). This compares to the $O(1)$ number of ancilla Majoranas [namely the 4 MZMs used to prepare the noisy state for injection, and the $O(1)$-length tails] required in the off-patch method from Section 4.4.2. Thus, for large $d$, utilizing fermionic aspects of the MSC in magic state preparation and storage becomes more efficient. (Note that it is not possible to use bosonic ancilla patches for measurement as we require odd-length boundaries for lattice surgery with triangle codes; see Table 3.)
@@
To consider the overheads associated with fermionic ancilla patches for logical measurement, we first discuss error processes that we must contend with in order for our procedures to be fault-tolerant. Measurement errors may make thin-width sections of code vulnerable to unobservable, low-weight error mechanisms, depending on the patch preparation and readout used. We illustrate this error through the example shown in Figure 16b (black ring). In this example, ancilla patch preparation starts by preparing blue bond operators in their $+1$ eigenstates, followed by the measurement of blue plaquettes, which all have random outcomes. A fermion-parity-odd string that terminates on red and green boundaries can flip a blue plaquette, as shown in Figure 16b. However, if such an error string occurs shortly after state preparation along blue bonds, measurement errors can make it unobservable: suppose, for example, a blue plaquette suffers an initial measurement error, and it is then flipped by the parity-odd string described above. Subsequent measurements of the blue plaquette cannot detect the error string because of the initial measurement error (see Ref. [Anyon_condensation_floquet_codes] for a discussion of analogous error mechanisms in the bosonic Color Code).
@@
6 Fault-Tolerant Gates on Small Majorana Surface Codes
@@
In Figure 6, we show how braiding two $Z_{3}$ twists in a quartet affects the logical operators $\bar{X}_{1}$, $\bar{Z}_{2}$ and $\bar{Z}_{1}$ for the two BT qubits. The effect on logical operator $\bar{X}_{2}$ is shown in Figure 10 of the main text.
@@ Appendix E Lattice Surgery for Large-Weight Logical MZM Bilinear Measurement
In Figure E, we illustrate the lattice surgery procedure for measuring the bilinear operator $i\Upsilon_{a}\Upsilon_{b}$, where these two logical MZMs have arbitrarily high weights. In the example, $\Upsilon_{a}$ is supported on a blue boundary and $\Upsilon_{b}$ on a red boundary. Similarly to in Section 4.3.1, we prepare a patch, supporting ancilla logical MZMs $\tilde{\Upsilon}_{a}$ and $\tilde{\Upsilon}_{b}$, in the state $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}=+1$. These logical MZMs will have a similar weight to $\Upsilon_{a}$ and $\Upsilon_{b}$ respectively, and will be supported on blue boundaries. We measure the extra plaquette operators shown in Figure 23b connecting the two patches. Note that the product of the eigenvalues of those plaquette operators connecting $\Upsilon_{a}$ and $\tilde{\Upsilon}_{a}$ is equal to the eigenvalue of $i\Upsilon_{a}\tilde{\Upsilon}_{a}$, and similarly for $i\Upsilon_{b}\tilde{\Upsilon}_{b}$. We then measure $i\tilde{\Upsilon}_{a}\tilde{\Upsilon}_{b}$ destructively (by measuring all blue bond operators) and perform/track a correction operation dependent on this measurement outcome (cf. Section 4.3.1).
@@ Appendix F Calculating the Correction Operation for Logical Fermion Parity Measurement
Here, we go through the details of Section 4.3.2, specifically showing how to determine the correction operation that needs to be tracked when measuring a logical fermion parity operator.
To determine this operation, we need to assess whether each logical bilinear operator has been preserved by the measurement-based procedure outlined. We deal with the example of $i\Upsilon_{1}\Upsilon_{3}$ specifically. We can use the final column of Table 2 to determine that after the lattice surgery measurements, this bilinear is mapped to $\lambda_{1}\lambda_{3}\bar{Z}_{B}$ (it is unchanged by the final measurement of $i\Upsilon_{4}\tilde{\Upsilon}_{4}$). The final step of the procedure from Section 4.3.2 is to measure all blue bond operators on the ancilla patch, from which we deduce the measurement outcome of operator $\tilde{\Gamma}_{4}$ to be $\eta_{1}$, $i\tilde{\Upsilon}_{1}\tilde{\Upsilon}_{2}\,\bar{X}_{B}$ to be $\eta_{2}$ and $i\tilde{\Upsilon}_{2}\tilde{\Upsilon}_{4}\,\bar{Z}_{B}$ to be $\eta_{3}$ (these measurements do not commute with $\bar{Z}_{B}$). The post-measurement state of the logical qubits can be obtained by acting three logical operators $R_{1}$, $R_{2}$ and $R_{3}$ on the pre-measurement state. This is equivalent to replacing projectors with logical braids in Section 4.3.1.
We now determine $R_{1}$, $R_{2}$ and $R_{3}$ for this example. Note the logical operators stabilizing the pre-measurement state are those that were just measured via lattice surgery: $\lambda_{j}i\Upsilon_{j}\tilde{\Upsilon}_{j}$ for $j=1,2,3,4$. We consider first the measurement of $\tilde{\Gamma}_{4}$ (it does not matter which we choose to consider first, since the measurements all commute). A logical braid that enacts the same transformation as this measurement is (see Section 4.3.1):
@@ Appendix G Arbitrary String Operator Measurement Details
Here we extend the procedures of Section 4.5 to measuring arbitrary products of fermionic and bosonic string operators. We assume we wish to measure a product of $k$ operators, labeled $S_{1},\ldots,S_{k}$ and colored either red or blue. We start by introducing $k$ ancilla patches, each hosting two logical MZMs, $\tilde{\Upsilon}_{j_{a}}$ and $\tilde{\Upsilon}_{j_{b}}$ for $j=1,\ldots,k$, and we prepare each patch in the $i\tilde{\Upsilon}_{j_{a}}\tilde{\Upsilon}_{j_{b}}=+1$ state. This also prepares the $+1$ eigenstate of operator $\tilde{\Gamma}=\prod_{j=1}^{k}i\tilde{\Upsilon}_{j_{a}}\tilde{\Upsilon}_{j_{b}}$. An example of this setup is shown in Figure LABEL:subfig:arb_mmt_app_1 for $k=2$. In this example, $S_{1}$ is a bosonic string operator encircling a pair of fermionic twists hosting logical MZMs on blue boundaries. Both $S_{1}$ and $S_{2}$ are red string operators. We then measure $i\tilde{\Upsilon}_{j_{b}}\tilde{\Upsilon}_{(j+1)_{a}}$ for all $j<k$, to entangle these ancilla patches, while preserving the state of $\tilde{\Gamma}$. To achieve this, we can bridge the gap between ancilla patches with other ancilla patches. In Figure LABEL:subfig:arb_mmt_app_1, we introduce a patch hosting $\tilde{\Upsilon}_{c}$ and $\tilde{\Upsilon}_{d}$. The measurement of $i\tilde{\Upsilon}_{1_{b}}\tilde{\Upsilon}_{2_{a}}$ is performed via lattice surgery, shown in Figure 24b (see also Section 4.3.1 for an explanation of this procedure). After this measurement, the middle patch hosting $\tilde{\Upsilon}_{c}$ and $\tilde{\Upsilon}_{d}$ is discarded.